In co, the J=0 →J=1 transition, absorption occurs at forequency 1-15x10" Hz. Moment of Inertia of the molecule is ? (b) find the Bondlength of the Co molecule.


1
Expert's answer
2021-06-21T11:40:35-0400

"\mu=\frac{m_1m_2}{m_1+m_2}=1.14\cdot 10^{-26}~kg,"

"\nu=\frac{1}{2\pi}\sqrt{\frac{k}{\mu}}\implies"

"k=4\mu \pi^2 \nu^2=0.45~\frac{MN}{kg},"

"I=\frac{h}{8\pi^2 k^3}=0.92\cdot 10^{-52}~kg\cdot m^2,"

"r=\sqrt{\frac{I}{\mu}}=8.98\cdot 10^{-14}~m."